I am solving Linear systems algebraically of equation's using substitution. The question is x+2y=2 and 5x-3y=-29

Answer by TimothyLamb(4379)

---
x+2y=2
5x-3y=-29
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix-method:
x = -4
y = 3
---
system matrix
1 2
5 -3
---
inverse of system matrix
0.23076923 0.15384615
0.38461538 -0.076923077
---
determinant of system matrix = -13
---
now solution by substitution-method:
---
x + 2y = 2
2y = 2 - x
y = 1 - (1/2)x
---
5x - 3y = -29
5x - 3(1 - (1/2)x) = -29
5x - 3 + (3/2)x = -29
(10/2)x + (3/2)x = -29 + 3
x = -26
x = -26(2/13)
---
x = -4
---
y = 1 - (1/2)x
y = 1 - (1/2)(-4)
y = 1 + 2
---
y = 3
---
solution by substitution-method:
x = -4
y = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php